10. You are the owner of a regional chain of and has persuaded you to try an experiment in your stores. Typically the sales in your et sales. She statistically equal but your daughter insists that if you play music in the stores it Win s stores where beginning to annoy you with her insistence so you go ahead conduct an you play slow music, fast music or no music. experiment but she does not explain. The manager is a psychology student your stores are will affect sales. She is experiment in 1S stores where Your daughter presents you with the findings of her at University of Baltimore and you ask her to explain. What does she tell your st assistant manager of one of your stores is studying busines does she provide? Your daughter did not conduct Tukey Kramer test but your assistant manager has a lot of initiative and she well take a qualified look at what is the most important hand. result with the results at ANOVA: Single Factor SUMMARY Groups Count Sum AverageVariance 286821.2 70209 75197 62824 14041.8 15039.4 326836.3 12564.8 638932.7 5 Slow Fast None ANOVA df MSF P-valueFcrit Source of Variation Between Groups Within Groups 2 7750316.6 18.5622958 0.0002124683.88529031 12 417530.067 15500633.2 5010360.8 20510994 14 Total
Answer #1

Null and Alternative hypothesis:

H_o:\mu_1=\mu_2=\mu_3

H_a:\textup{At least one mean is not equal.}

As we can interpret from the table:

Fstat = 18.562

And Fcrit = 3.885

As Fstat =18.562 > Fcrit = 3.885 , we reject the null hypothesis.

There is enough evidence to claim that not all means are equal.

Tukey Kramers Test.

\alpha = 0.05

Number of treatment, k = 3

Number of value in each treatment, n= 5

Total value, N = 15

MSE = 417530.067

Critical value, q_{\alpha /2, k, N-k}=q_{0.025, 3, 12} = 3.775

Margin of error, E =

E =q_{0.025, 3, 12}*\sqrt{\frac{MSE}{n}} = 3.775*\sqrt{\frac{417530.067}{5}}= 1090.877

Difference (\mu_i-\mu_j)\pm E Interval Differs or not
\mu_1-\mu_2 (14041.8-15039.4) \pm 1090.877 (-2088.477, 93.277) No
\mu_1-\mu_3 (14041.8-12564.8) \pm 1090.877 (386.123, 2567.877) Yes
\mu_2-\mu_3 (15039.4-12564.8) \pm 1090.877 (1383.723, 3565.477) Yes

As we can see that there is if we play music (slow and fast) the sales is more as compared to the sale when no music is played.